Thursday, 20 August

Other Events

12:15  Springer Booth, Exhibition Area

Presentation of Olli Lehto’s New Book “Mathematics Without Borders, A History of the International Mathematical Union”

Reception on the occasion of the publication of this book, organized by Springer-Verlag and IMU

\underline{18:30 Urania}

Opening of the Exhibitions Mathematik zum Anfassen — Hands–on Mathematics, Mathematical Stone Sculptures, Mathematics and Ceramics Works by High–School Students

\underline{19:00 Urania}

Opening of the VideoMath Festival

\underline{19:30 Urania}

Frank Natterer, Münster, Germany The Mathematics of Imaging in Medicine and Technology

ICM’98 – Program  1